# SIC-POVMs and orders of real quadratic fields

### v1

@@ 6. Theorems on class monoids of real quadratic fields
In this section, we prove rigorous results about the behavior of $\operatorname{Clm}(\mathcal{O})$ for general orders. These results have conditional consequences for SICs via 5.4.
6.1. Growth of the number of SICs in dimension $d$
@@
Going beyond Zauner’s conjecture, one would like to know not only that SICs exist in every dimension $d$, but also how many there are. The connection between SICs and class numbers gives us a precise prediction for the number of SICs up to equivalence (excluding any SICs that are not Weyl–Heisenberg covariant, to which our conjectures are agnostic). In this section, we use known results on the growth of class numbers to derive conditional results on the number of SICs as $d\to\infty$.
@@
Set $\Delta=\Delta_{d}=(d+1)(d-3)$ and $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta})\right|$; we are trying to determine when $s(d)=1$. By 5.4,
@@
where $\Delta=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental.
If $f>1$, then $s(d)>1$.
If $f=1$, then $\Delta=\Delta_{0}$ is fundamental. Write $\Delta=(d-1)^{2}-4=n^{2}-4$, where $n:=d-1$. Byeon, Kim, and Lee solved the class number 1 problem for real quadratic fields with fundamental discriminant of the form $n^{2}-4$, a result previously known as Mollin’s conjecture [13, Thm 1.2]; whenever $n>21$ (so $d>22$), $s(d)=h_{\Delta}>1$. The list of dimensions where $s(d)=1$ follows from a finite calculation of class numbers. ∎
@@
The corollary follows from 6.1. ∎
The second growth result is an asymptotic formula for the size of the ray class monoid of $\mathcal{O}_{\!\Delta_{d}}$, which is conjectured to be the equal to the number of extended Clifford orbits of SICs in dimension $d$. The implied constant in this formula is ineffective because it ultimately relies on a Siegel bound involving a possible Siegel zero of a Dirichlet $L$-function. We are grateful to Steven Flammia for independently finding numerical evidence for this asymptotic [23].
@@
As $d\to\infty$, and $\Delta_{d}=(d+1)(d-3)$ the size of the class monoid of $\mathcal{O}_{\!\Delta_{d}}$ obeys the asymptotic formula
@@
The proof uses the Brauer–Siegel theorem, which is an (ineffective) estimate on the growth of the product of the class number and the regulator. The regulator is always negligibly small in our regime. The subtlety of the proof is entirely in accounting for the class numbers of non-maximal orders, which are not included in the Brauer–Siegel theorem.
As in the proof of 6.1, set $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$. By 5.3,
$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}$ (6.3)
where $\Delta=\Delta_{d}=(d+1)(d-3)=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental. Also, $h_{(f^{\prime})^{2}\Delta_{0}}\leq h_{\Delta}$ for all $f^{\prime}\mid f$. It follows that
$h_{\Delta}\leq s(d)\leq\sigma_{0}(f)h_{\Delta},$ (6.4)
where $\sigma_{0}(f)$ is the number of positive integer divisors of $f$. Moreover,
$\log\sigma_{0}(f)\ll\frac{\log f}{\log\log f}\ll\frac{\log d}{\log\log d}$ (6.5)
by [34, Thm. 2.11]. Thus,
$\log s(d)=\log h_{\Delta}+O\!\left(\frac{\log d}{\log\log d}\right).$ (6.6)
We then use the formula for the class number of an order given as [35, Thm. 12.12]:
$h_{\Delta}=\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}.$ (6.7)
We will estimate the size of each factor in this formula. For any real nonsquare discriminant $\Delta^{\prime}$, let $\varepsilon_{\!\Delta^{\prime}}$ be the smallest unit of $\mathcal{O}_{\!\Delta^{\prime}}$ such that $\varepsilon_{\!\Delta^{\prime}}>1$, and let $\varepsilon_{\!\Delta^{\prime}}^{+}$ be the smallest such totally positive unit. The Brauer–Siegel theorem [12]—specifically, Siegel’s part, concerning real quadratic fields [40]— gives the asymptotic formula
$\frac{\log\!\left(h_{\!\Delta_{0}}\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))\right)}{\log{\Delta_{0}}}\sim 1\mbox{ as }\Delta_{0}\to\infty.$ (6.8)
Moreover,
$[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]=\frac{\log\varepsilon_{\!\Delta}}{\log\varepsilon_{\!\Delta_{0}}}=\frac{\log\varepsilon_{\!\Delta}}{\log\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))}.$ (6.9)
Thus, (6.8) may be rewritten as
$\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\log\varepsilon_{\Delta}\right)\sim\log{\Delta_{0}}\mbox{ as }\Delta_{0}\to\infty.$ (6.10)
Fix some $\epsilon>0$. Then, there exists some natural number $N$ such that, whenever $\Delta_{0}>N$,
$(1-\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}<\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)<(1+\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}.$ (6.11)
For $\Delta_{0}\leq N$, there is some constant $B$ such that
$B-\log\log\varepsilon_{\Delta}\leq\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)\leq B-\log\log\varepsilon_{\Delta}.$ (6.12)
We may bound $\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ above by $f^{2}$ and below by
$\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ $=f^{2}\prod_{\mathfrak{p}|f}\left(1-\frac{1}{\operatorname{Nm}(\mathfrak{p})}\right)\geq f^{2}\prod_{p|f}\left(1-\frac{1}{p}\right)^{2}=\varphi(f)^{2},$ (6.13)
where the first product is over prime ideals of $\mathcal{O}_{\!\Delta_{0}}$ and the second is over rational primes. Since $\sqrt{\Delta}=f\Delta_{0}\equiv 0\,\left(\operatorname{mod}\,f\mathcal{O}_{\!\Delta_{0}}\right)$, the multiplicative group $\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\cong\left(\mathbb{Z}/f\mathbb{Z}\right)^{\times}$, so $\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|=\varphi(f)$. Thus,
$\varphi(f)\leq\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq\frac{f^{2}}{\varphi(f)}.$ (6.14)
Using (on both sides) the inequality $\varphi(f)>\frac{f}{\ell(f)}$ where $\ell(f)$ is a real-valued function satisfying $\ell(n)>1$ for all $n$ and $\ell(n)=e^{\gamma}\log\log n$ as $n\to\infty$ [34, Thm. 2.9], we have
$\frac{f}{\ell(f)}<\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq f\ell(f),$ (6.15)
which after taking logarithms becomes
$\log f-\log\ell(f)<\log\!\left(\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\right)<\log f+\log\ell(f).$ (6.16)
Using (6.7) and combining (6.11) and (6.16), for $\Delta_{0}>N$, we obtain
$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<\epsilon\log\sqrt{\Delta}_{0}+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f).$ (6.17)
On the left-hand side, $\log\sqrt{\Delta}_{0}+\log f=\log\sqrt{\Delta}\sim\log d$ as $d\to\infty$. On the right-hand side, as $d\to\infty$, $\log\sqrt{\Delta_{0}}<\log\sqrt{\Delta}\ll\log d$, while
$\log\log\varepsilon_{\Delta}\leq\log\log\varepsilon_{\Delta}^{+}=\log\log\frac{d-1+\sqrt{\Delta}}{2}\ll\log\log d,$ (6.18)
and $\log\ell(f)\ll\log\log\log d$. Thus,
$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.19)
In the case when $\Delta_{0}\leq N$, combining (6.11) and (6.17), we instead obtain
$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<B+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f),$ (6.20)
which by a similar argument also implies
$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.21)
Combining (6.19) and (6.21) with (6.6) proves the theorem in all cases. ∎
@@
Follows from 6.3. ∎
7. Conjectures on counting Galois multiplets of Weyl–Heisenberg SICs
Recall that 4.10 implies that there are a finite number of well-defined Galois multiplets of SICs in dimension $d$, each consisting of a finite number of geometric equivalence classes; we will now refine this prediction to a precise count in every dimension. Our Multiplet Count Conjecture (7.1) predicts the number of multiplets in dimension $d$. The Order-to-Multiplet Conjecture (7.2) predicts the number of geometric equivalence classes in each multiplet.
7.1. Statement of the Multiplet Count Conjecture
[…]
Conjecture 7.1 (Multiplet Count Conjecture).
Fix a positive integer $d\neq 3$. Let $\Delta=\Delta_{d}=(d+1)(d-3)$ and $K=\mathbb{Q}(\sqrt{\Delta})$. The number of Galois multiplets of Weyl-Heisenberg SICs in dimension $d$ equals the number of quadratic orders $\mathcal{O}^{\prime}$ with $\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$. That is,
where $\sigma_{0}(f)$ denotes the number of positive divisors of $f$.)
The content of 7.1 is that the multiplets of Weyl–Heisenberg SICs are in one-to-one correspondence with orders between $\mathcal{O}$ and $\mathcal{O}_{K}$. The fact that the number of such orders is simply $\sigma_{0}(f)$ uses the fact that an order in a quadratic field is determined uniquely by its discriminant, which can be $(f^{\prime})^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$ and a positive integer $f^{\prime}$. The order $\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ contains $\mathcal{O}_{f^{2}\Delta_{0}}=\mathcal{O}_{\!\Delta}$ if and only if $f^{\prime}\mid f$. (These facts are not true for general number fields.)
7.2. Statement of the Extended Order-to-Multiplet Conjecture
We next formulate a conjecture positing a correspondence between orders and multiplets with certain properties. The number of geometric equivalence classes $[S]$ inside a Galois multiplet $[[S]]$ should be equal to the class number of the corresponding orders, and the fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ associated to different multiplets are related in a natural way. (Recall that we have already shown these fields to be are well-defined on $\operatorname{PEC}(d)$-equivalence classes, and they are well-defined on Galois super-equivalence classes under the assumption of 4.10.)
Conjecture 7.2 (Extended Order-to-Multiplet Conjecture).
Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ for a fundamental discrimiant $\Delta_{0}$ and some positive integer $f$. Then there is a bijection $\mathcal{M}$,
{positive divisors of $\{\mbox{positive divisors of $f$}\}\xrightarrow{\mathcal{M}}\{[[S]]:S\mbox{ a Weyl--Heisenberg line-SIC in }\mathbb{C}^{d}\}$ (7.3)
having the following properties:

### v3

@@ 7. Theorems on sizes of class monoids of real quadratic fields
In this section, we state rigorous results about the behavior of $\left|\operatorname{Clm}(\mathcal{O})\right|$ for orders with discriminants $\Delta_{d}=(d+1)(d-3)$. These results have conditional consequences for SICs via 5.4.
@@
Going beyond Zauner’s conjecture, one would like to know not only that SICs exist in every dimension $d$, but also how many there are. The connection between SICs and class numbers gives us a precise prediction for the number of SICs up to equivalence (excluding any SICs that are not Weyl–Heisenberg covariant). In this section, we use known results on the growth of class numbers to derive conditional results on the number of SICs as $d\to\infty$.
@@
Set $\Delta=\Delta_{d}=(d+1)(d-3)$ and $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta})\right|$; we are trying to determine when $s(d)=1$. By 5.3,
@@
where $\Delta=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental. If $f>1$, then $s(d)>1$.
If $f=1$, then $\Delta=\Delta_{0}$ is fundamental. Write $\Delta=(d-1)^{2}-4=n^{2}-4$, where $n:=d-1$. Byeon, Kim, and Lee solved the class number 1 problem for real quadratic fields with fundamental discriminant of the form $n^{2}-4$, a result previously known as Mollin’s conjecture [11, Thm 1.2]; whenever $n>21$ (so $d>22$), $s(d)=h_{\Delta}>1$. The list of dimensions where $s(d)=1$ follows from a finite calculation of class numbers. ∎
@@
The corollary follows from 7.1 combined with 5.4. ∎
The second growth result is an asymptotic formula for the size of the ray class monoid of $\mathcal{O}_{\!\Delta_{d}}$. The proof uses the Brauer–Siegel theorem, which is an (ineffective) estimate of the growth of the product of the class number and the regulator for maximal orders. The presented result for orders must account for the class numbers of non-maximal orders. The conductor $f=f_{\Delta_{d}}$ is an extra degree of freedom not present in the Brauer–Siegel theorem.
@@
As $d\to\infty$, and $\Delta_{d}=(d+1)(d-3)$, the size of the class monoid of $\mathcal{O}_{\!\Delta_{d}}$ obeys the asymptotic formula
@@
This is proved in [36], using results from [35]. ∎
@@
The corollary follows from 7.3 combined with the conjecture. ∎
8. Theorems on strictness or nonstrictness of inclusions of ray class fields of orders
In [36] we prove unconditional results determining exactly when the fields $\tilde{E}_{d,f^{\prime}}$ introduced in 6.3 can coincide, while fixing $d$ and varying $f^{\prime}$.
Theorem 8.1.
The set of $d\geq 4$ such that, for $\Delta=(d+1)(d-3)$ and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$, there exist distinct orders $\mathcal{O}\neq\mathcal{O}^{\prime}$ with
$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (8.1)
having $H_{d\mathcal{O},\Sigma}^{\mathcal{O}}=H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}$ for some $\Sigma$ (equivalently, all $\Sigma$) is exactly the set of all $d$ such that the squarefree part of $\Delta$ (called the radical $\operatorname{rad}(d)$) has rad(Δ)1(mod $\operatorname{rad}(\Delta)\equiv 1\,\left(\operatorname{mod}\,8\right)$.
The following theorem classifies all pairs where coincidence may occur.
Theorem 8.2.
Let $d\geq 4$. Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for $\Delta_{0}$ a fundamental discriminant, and suppose $f^{\prime},f^{\prime\prime}$ are positive integers satisfying $f^{\prime}\mid f$, $f^{\prime\prime}\mid f$ and $f^{\prime\prime}<f^{\prime}$. The following are equivalent: